OpenStudy (anonymous):

Can I get some people to help me with finding Areas of circles? Five questions.

OpenStudy (anonymous):

yes i will help

OpenStudy (black_ninja123):

I will help also

OpenStudy (anonymous):

Okay Ima going to start the quiz and I will post the q then we can find the a together.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the exact area of a circle with a radius of 12 feet?

OpenStudy (anonymous):

24 (pi) square feet?

OpenStudy (black_ninja123):

3.14*12^2

OpenStudy (anonymous):

What does ^ mean?

OpenStudy (black_ninja123):

to the power of

OpenStudy (anonymous):

452.39

OpenStudy (anonymous):

I'm looking for the exact not the approximate biloxid. So the answer choices are 24 (Pi) square feet 36 (pi) square feet 121 (pi) square feet 144 (pi) square feet

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

Its 24 right @Black_ninja123

OpenStudy (black_ninja123):

It will be 144pi because 12 to the power of 2 is 144

OpenStudy (anonymous):

Sheezus okay I get it know here the next q.

OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 18 feet? 36 square feet 81 square feet 324 square feet 1296 square feet

OpenStudy (anonymous):

With (pi) after eah chouce.

OpenStudy (black_ninja123):

9 to the power of 2= 81

OpenStudy (black_ninja123):

81pi is the answer

OpenStudy (black_ninja123):

Do you get how to do it now?

OpenStudy (anonymous):

So I take the radious of the circle and essently ^2 it to get my answer?

OpenStudy (black_ninja123):

yes because the formula is pi * r^2

OpenStudy (anonymous):

Okay but the next questions are different so I still need you help. What is the approximate area of a circle with a radius of 28 centimeters? Use (pi) = 3.14. 87.9 square centimeters 175.8 square centimeters 2461.8 square centimeters 4923.5 square centimeters

OpenStudy (black_ninja123):

You do this by yourself and tell me what you get

OpenStudy (black_ninja123):

Its the same thing but this time you have to multiply 3.14

OpenStudy (anonymous):

3.14 * 2 * 28

OpenStudy (anonymous):

I latterly don't know because I've been busy with science but my best guess is: 28 +28 = 56^2? could you give me the formula?

OpenStudy (anonymous):

\[\pi \times radius ^{2}\]

OpenStudy (black_ninja123):

3.14*14^2=?

OpenStudy (anonymous):

I'm looking for the approximate not exact @isthatguy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

How do I find the approximate are and how do I find the exact area?

OpenStudy (black_ninja123):

for exact you dont use 3.14 and for approximate you use 3.14

OpenStudy (anonymous):

well multiply 3.14 by 28^2 or 3.14 by (28x28) to get your approximate answer

OpenStudy (anonymous):

Ah okay thanks let me try to solve the rest of the quiz on my own but if I do bad I'll come back to see what I messed up on.

OpenStudy (black_ninja123):

Ok

OpenStudy (anonymous):

@itsthatguy yeah thats what i was going for.

OpenStudy (anonymous):

well, that's cool

OpenStudy (anonymous):

Thanks guys got a 5/5 I wish I could give you all medals C: except you. You know who you are... XD